I need help on steps for solving a system of equations in substituting,elimination,and graphing method for problem
3x-y=7
Y++2x=8
Thank you in advance
I worked the problem
3x=(2x-8)=7
3x-2x+8=7
X+8=7x=-1
-8-2=-10
(-1,-10)

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

3x-y=7
Y++2x=8 (one + is sufficient)
Thank you in advance
I worked the problem
3x=(2x-8)=7 *** I'll assume the 1st = is a typo, and should be -
3x-2x+8=7 *** OK so far
X+8=7x=-1 *** again, too many equal signs
-8-2=-10
(-1,-10)
------------------
Start at the beginning.
3x-y=7
Y+2x=8 --> y = -2x + 8
sub for y in the 1st eqn
3x - (-2x+8) = 7
3x + 2x - 8 = 7
5x = 15
x = 3
---------
sub for x in any equation to find y
Y++2x=8
y + 2*3 = 8
y = 2
====================
You can do these, you're just careless and make mistakes.

Answer by ewatrrr(24785)

3x-y=7 0r y =3x-7
Y + 2x=8
(3x-7) + 2x = 8
5x = 15
x = 3 and y = 2 9-7
(3,2) the solution to this System
..........
Elimination
3x - y = 7
2x + y = 8
5x = 15, x = 3 and y = 2
.........
Graphing
y = 3x-7 P(0,-7) and (2,-1) on this line.. Plot points and connect with Line
y = -2x + 8 P(0,8) and P(4,0) on this line....Plot points and connect with Line
Intersection: P(3,2)